CASE 4-9 o Estimated Medicare-Severity Diagnosis-Related Group Payments Agministration wants so know teM ment for several MS-DRGs. The grouper an your s rs. The haspitas base rate is $7,320.50 Th M5 09Os Thot admiistraisn is concerned sbest ie Table 4-17 lse Table46 to enter the relative weight fram Toble 4-7 and celculate the essmate payer en tates vailable on the studest companion wabaste tor your ine MS-DRGs with Relative Weight and Estimated Payment MS-DRG MS-DRG Title 190 Estimated Payment Relative Weight 193 231 281 Chronic obstructive pulmonary disease with MCC Simple pneumonia and pleurisy with MCC Coronary bypass with PTCA with MCC Acute myocardial infarction discharged alive with CC Hypertension with MCC 304 334 374 389 472 509 Rectal resection without MCC/CC Digestive malignancy with MCC Gl obstruction with CC Cervical spinal fusion with CC Arthroscopy

media%2Fe66%2Fe66cebe5-f5ae-4294-aa4e-05
Answer #1

The MS-DRGs identify the group of patients with similar clinical problems and are expected to consume similar amounts of hospital resources. The grouping is done based on factors such as patient diagnoses or surgery. The relative weight of each MS-DRG compares its costliness to the average Medicare case.

-Base payment rate comprised of the amount that divided into labour and non-labour shares. The labour-related share is adjusted by a wage index applicable to the hospital’s geographical location.

The system uses DRG-specific weights to calculate a payment above or below the fixed payment, the base payment rate (operating and capital). The base payment rate reflects the cost (labour and non-labour) to deliver care to a patient for an average Medicare hospitalization.

Add on is being made to the IPPS payment for hospitals that serve a disproportionate share of low-income patients as DSH adjustment and approved teaching hospitals that incur indirect costs of medical education as IME adjustment.

Add on % = 2%

Hospital base rate: $7320.50

Base rate adjusted for geographic factor x DRG weight = DRG

Estimated payment = Base rate x MS-DRG relative Weight + add on

MS- DRGs Estimated payment

1. 190                   7320.50 x 1.1743 + 2 = 8598.5

2. 193                   7320.50 x 1.4491 + 2 =10610.1

3. 231                   7320.50 x 7.7247 + 2 =56550.7

4.281 7320.50 x 1.0247 + 2 =7503.3

5. 304                   7320.50 x 1.0016 + 2 =7334.2

6. 334                   7320.50 x 1.6032 + 2 =11738.2

7. 374                   7320.50 x 2.0182 + 2 =14776.2

8. 389                   7320.50 x 0.8717 + 2 =6383.3

9. 472                   7320.50 x 2.9166 + 2 =21352.9

10. 509                 7320.50 x 1.5499 + 2 =11348.04